At the beginning of the day, the cash register contains $150. At the end of the day, it contains $1114.80. If the sales tax rate is 7.2%, how much of the total is sales tax?
Answer by checkley71(8403)

1114.80-150.00=964.80 thus x+.072x=964.80 or 1.072x=964.80 or x=900 thus the sales were $900.00 with an added sales tax of 7.2%
900+.072*900+150=1114.8 or 900+64.8+150=1114.8 or 1114.8=1114.8